"A taxi driver would like to earn $5,000 per year in supplemental income from an inheritance of $80,000. If the driver invests $20,000 in a mutual fund paying 8%, what rate must he earn on the remainder to achieve his goal?"
I'm just having trouble putting the information given into an equation to be solved.

Answer by richwmiller(17219)

a+b=80000
a=20000
b=60000
.08*20000+x*60000=5000
.08*20+x*60=5
.08*4+x*12=1
.32+12x=1
12x=.68
x=0.0566667=5.66667%
check
.08*20000+.0566667*60000=5000.002
ok